two bodies of different masses ma and mb are dropped from two different heights a and b.the ratio of times taken by the two bodies to drop through these distance is:

Grade:12

2 Answers

Sudheesh Singanamalla
13 years ago

mass of 'A' is Ma

mass of 'B' is Mb

both of them undergo an accelaration of 'g' as it is falling from a height

initial velocity of both is zero as it is dropped from rest.

From 2nd equation of motion.

S = 1/2 a t2                      [ since u = 0 , it is not taken. ]

a = 1/2g t12

b = 1/2g t22

a/b = t12 / t22

t1 / t2 = √a/b

 

this the the ratio of times taken by Ma and Mb

 

Please approve

ankit singh
3 years ago
for First object we get :

h1= gT1²/2
⇒T1=√2h1/g

For second object we get h2=gT2²/2
⇒T2=√2h2/g
T1/T2 =√h1/h2
But here h1= a and h2 =b
so, T1/T2=√a/b